Results from the CDF Detector at Fermilab

1. p p Physics

2. The CDF Detector

3. Electroweak Results

4. QCD Physics

5. Heavy Flavour Physics

6. The Next Run

I. p p Physics at the Tevatron Collider

The Standard Model has been enormously successful

• predicts the structure of the electroweak interaction

• describes both low and high-energy behaviour

• properties of the Z and W bosons

Study of high energy proton-antiproton collisions

• probe electroweak sector

• test QCD

• study heavy quark production

• search for new physics

The CDF Collaboration

The CDF collaboration is a group of 20 institutions

• studying p p interactions at √s = 1.8 TeV.

• Focus is on high-pT physics

Data collected

• Engineering run in 1987 – collected 30 pb−1 of “minimum bias” data

– 22 W → eνe candidates

– measured jet and inclusive particle pT spectrum

• Had first major run in 1988-89 – collected 4.4 pb−1 with full L1/L2/L3 trigger – ∼ 3000 W → eνe candidates

– ∼ 15 publications with several more in progress

The CDF Detector

• CDF is a general purpose, solenoidal detector

• Solenoid magnet of radius ∼ 1.5 m and length 5 m generating a 1.4 T magnetic field.

• Tracking chambers inside magnetic field provide particle

momentum resolution of

\frac{σp}{p} ∼ 0.0011p

• Calorimeters surround the interaction point

• Used for electron and muon identification

• Detection of jets

• Measurement of underlying event

• Drift chambers outside of calorimeter for muon ID

Lepton ID

Electron ID provided by

• Transverse shower profile

• Longitudinal energy deposition (had/em)

• Shower profile measurement using PWC’s imbedded in EM • Shower-track position match

• Calorimeter Energy-track momentum match ⇒ S/N ∼ 5 − 10 for inclusive e±

Muon ID uses drift chambers outside of calorimeter

• Matching of muon track stub and track in CTC

• Minimum ionizing energy deposition in calorimeter

Neutrino’s identified by presence of missing E_{T} (E̸_{T})

• Provided by almost hermetic calorimetry over |η| < 4.2

Data Collection

Focus is on high-p_{T} physics

• total cross section is ∼ 50 mb

• “interesting physics” cross section is ∼ 10 − 100 nb

• Decision to keep an event made by a 3-level trigger system: – L1: 90% of events are thrown out after the first 3 μs – L2: 99.9 % are thrown out after 100 μs

– L3: 70% removed using a software algorithm

• Collected 4.4 pb^{−1} of data in 1988-89

– (5 × 106 events)

ElectroWeak Physics

Properties of the W and Z bosons:

• First measurement was mass and width of the Z – important test of detector systems

– used best measured e± and μ± events

– calibrated tracking chamber in situ MZ = 90.9 ± 0.3(stat. + syst ) ± 0.2(scale ) GeV/c2

• Measured the mass and width of the W – Indirectly measure ν PT (but not PL)

– Use the transverse mass

M{eν \atop T} ≡ √2ET E̸T (1 − cos ϕeν)

Use of E̸T requires good understanding of calorimeter – Reduced these by considering events with no jets

– 1130 eνe and 592 μνμ candidates MW = 79.91 ± 0.39 GeV/c2

Results in a measurement of sin2 θW = 0.231 ± 0.008

Studies of Electroweak Boson Production

• W and Z Cross section ratio

\frac{σ(W → eνe)}{σ(Z → ee)} = \frac{σ(pp → W X)}{σ(pp → Z)} \frac{Γ(W → eν)}{Γ(Z → ee)} \frac{Γ(W)}{Γ(Z)}

gives another test of the standard model

• Selected sub-sample of W and Z events – 1828 W → eνe and 193 Z → e+ e− events

– \frac{σ(W →eνe)}{σ(Z→ee)} = 10.2 ± 0.8 ± 0.4

• Using LEP Γ(Z) measurement, can determine Γ(W) = 2.19 ± 0.20 GeV

• If assume Nν = 3, then Mtop> 35 GeV/c2 at 95% CL.

Weak Boson Production Asymmetries

The SM predicts an asymmetry in Z production

• results from the form of the neutral current

J{NC \atop μ} = J{I3 \atop μ} − sin2 θW J{EM \atop μ}

• expect e+ to be aligned with q̄

• observed asymmetry is AF B = [5.3 ± 5.9 (stat ) ± 0.4 (sys)] %

• Implies that sin2 θW = 0.228{+0.017 \atop −0.015} (stat ) ± 0.002(sys)

Expect W^{−} asymmetry in the decay mode e^{−}ν_{e}

• e− is aligned with d in initial state

• Sensitive to relative u and d structure functions

Perturbative QCD Physics There has been extensive studies of the inclusive jet data and the multijet events

• Inclusive jet cross section

• Sets a compositeness limit of ΛC > 1400 GeV/c2 (95% C.L. )

Kinematics of multi-jet events test QCD predictions

• Di-jet event mass spectra sensitive to higher-order QCD corrections

• Three-jet event kinematics also provide QCD test

Exotic Particle Searches

Searched for exotic phenomena:

• Extensions of standard model (W ′’s and Z′’s) – Search for analomolous high mass events

– See no excess of events in lepton + neutrino channels – Set limit of

MW ′ > 520 GeV/c2 (95%C.L.)

– See no evidence in either e+ e− or μ+ μ− channels – set limit of

MZ′ > 380 GeV/c2 (95%C.L.)

Heavy Flavour Physics

Tevatron Collider is a copious source of heavy quarks

CDF collaboration completing several studies

• b-quark cross section

• B∘ − B̄∘ Mixing

• Reconstruct b-quark exclusive decay modes

• Search for rare decays

• Search for heavier quarks (top and b′)

Bottom Quark Production at √s = 1800 GeV

CDF inclusive e^{±} sample used to estimate σ_{b} The event sample selected by requiring:

• A well-identified electron candidate – EM cluster with shower profiles consistent with electron – Shower leakage < 4% into hadronic calorimeter

– Only 1 charged particle pointing to cluster

– 0.75 < E/p < 1.4

• Remove conversion electron candidates – veto events with 2nd track consistent with γ → e+ e− – Remaining conversion background 20 ± 5%

Remaining charge hadron background is estimated at 15 ± 15% Sources of “prompt” electrons are:

• W candidates – Vetoed by requiring E̸T < 8√ET

• Z∘ candidates

– Vetoed events with 2nd e± candidate, m(e+ e−) > 80

GeV/c2

• bottom and charm production

Uncertainties in σ_{b} are: Resulting cross sections are:

B^{∘}−B̄^{∘} Mixing

Mixing in the B^{∘}−B̄^{∘} system is predicted to arise from the familiar “box-diagrams:”

It depends on

• the b quark lifetime,

• the CKM matrix elements, and

• the mass of the top quark.

Mixing is expressed in terms of

χ ≡ \frac{P rob(B̄∘ → B∘)}{P rob(B̄∘ → B̄∘) + P rob(B∘ → B∘)}

Cleanest signature comes from events where both B mesons decay semi-leptonically

• Can in principle have large rate of events

• Measure rate of like-sign (LS) to opposite-sign (OS) di-

leptons

R = \frac{N(l+ l+ ) + N(l−l−)}{N(l + l−)}

Several complications

• Cannot distinguish between B∘ s and B∘ d mesons

⇒ measure χ̄ ≡ f_{s}χ_{s} + f_{d}χ_{d}

where fd and fs are relative fractions of B∘ d and B∘ s mesons produced.

• Have to account for several potential backgrounds – Sequential decays B̄∘ → cX with c → l+ νlX

– Sources of “fake” lepton candidates

– Other sources of real di-lepton events

• Acceptance depends on b production and decay kinematics – Rely on MC when extracting χ̄ from R

B^{∘}−B̄^{∘} Mixing in Electron-Muon Events

CDF has measured mixing using the eμ final state

• Rate is double the di-muon or di-electron channel

• No backgrounds from Drell-Yan di-lepton production

Event Selection

• Event passed the eμ trigger requirements:

– electron candidate with ET > 5 GeV

– muon candidate with PT > 3 GeV/c

• Require a well-identified electron candidate – EM cluster with < 4% hadronic leakage

– transverse shower profiles consistent with e± shower – track extrapolated to shower centroid with E/p < 1.5.

• Require a muon candidate

– well-measured CTC track consistent with coming from the primary vertex

– good match between CTC track and μ chamber stub – minimum-ionizing track in calorimeter

Background Removal

• Remove W bosons by requiring E̸T < 15 GeV • Remove conversions identified by either

– finding a second track making γ → e+ e− – no associated VTPC track with electron

There are no backgrounds from

• Drell-Yan or Z∘ production

• J/Ψ or Υ resonances Results in 429 like-sign and 911 opposite-sign events.

Sequential Decays

• Half of the sequential decays come from the same B

b → μ−cX → μ−e+ X′

– These can be removed by requiring m(eμ) > 5 GeV – preferentially removes events where both b’s are nearby

• Remaining sequential background subtracted using the MC

Results in 346 like-sign and 554 opposite-sign events.

Background from fake lepton candidates

• Estimated probability that track is labelled μ candidate • Multiplied by number of such tracks in eμ data

• Results in background of

20 ± 10%

The background contributes equally to the LS and OS sample so that

R = \frac{346 − 90}{554 − 90} = 0.552 ± 0.049(stat )_{−0.048}^{+0.039} (bkg )

Charm and b sequential decays small background

• Estimated using ISAJET Monte Carlo

• Took into account higher-order production diagrams

Can extract χ̄ using

R = \frac{2χ̄ (1 − χ̄) + [(1 − χ̄)^{2} + χ̄^{2}] N_{s}/N_{f}}{[(1 − χ̄)^{2} + χ̄^{2}] + 2χ̄ (1 − χ̄) N_{s}/N_{f} + N_{c}/N_{f}}

• Ns = number of secondary b-quark decays

• Nf = number of direct b-quark decays

• Nc = number of c-quark decays

Resulting value for mixing is χ̄ = 0.176 ± 0.028(stat ) ± 0.025(MC) ± 0.032(MC)

Upper Limits on B → μ^{+} μ^{−}

CDF has searched for the flavour-changing decay

B → μ^{+} μ^{−}

• allowed in Standard Model at 10−9 − 10−8 level

• limits from e+ e− experiments ∼ 5 × 10−5

Select dimuon events using standard muon identification cuts

• Search for B signal in m(μ+ μ−) spectrum

• Resulting limits are:

BR(B → μ+ μ−) < 3.2 × 10−6 (Preliminary )

Search for Exclusive B Decays

Studied J/Ψ sample

• Very clean sample (S/N > 10)

• Expect to have large contribution from B decay

• Have searched for evidence of exclusive B decays in – B → J/ΨK

– B → J/ΨK∗(890)

Selected J/ΨK candidate events by requiring:

• J/Ψ PT > 8 GeV/c

• charged track with PT > 3.0 GeV/c in 60∘ cone around J/Ψ

Performed similar search in J/ΨK^{∗} mode:

• Consider 3 highest PT tracks in event only in 60∘ cone around J/Ψ

• Required 0.846 < m(Kπ) < 0.946 GeV/c2

Combined signal results in 35 ± 9 events

Analyses underway to

• measure B meson cross section • estimate the B meson lifetime

The Top Quark Search

Top quark remains an important standard model test

• Theory riddled with inconsistencies if top quark doesn’t exist with a mass < ∼ 180 − 200 GeV/c2

• CDF is currently the most sensitive experiment the next few

years

Search looks for several classes of events

• lepton + jets final states

• Dilepton final states

Most sensitive technique combines electron and muon channels

• electron or muon + ≥ 2 jets + b− → μX

• electron + muon (PT > 15 GeV/c)

• dielectrons and dimuons – Removed Z’s in e+ e− and μ+ μ− channel

Found one candidate event (in eμ channel)

• has a forward μ candidate (from B?)

• not typical of primary background (Z → τ+ τ−)

• Assuming candidate is real t t, set mass limit of 28 < Mtop < 91 GeV/c2 (at 95% CL)

A Look into the Future

CDF will acquire ∼ 5 times more data in next two years

• Should improve W mass measurment by at least 2-3

• Top search sensitive to a mass as large as 130-140 GeV/c2 • Expect ×10 more b-quark candidates

• Increased statistics will benefit other studies

Future runs at the Tevatron collider being planned

• Run in 1994 will obtain another factor of 5 in sensitivity • Will either have

1. discovered the top quark, or

2. found a clear contradiction with the Standard Model